A 1 m telescope can collect a given amount of light in 1 hour. Under the same observing conditions, how much time would be required for a 7 m telescope to perform the same task?
Answer #1

Given Data

distance d1 = 1 m

time T 1= 1 hour = 60 min

distance d2 = 7 m

time T 2 = ?

Solution :-

A telescope collects light in a rate proportional to the area of its primary mirror which is proportional to the diameter squared.

T1/T2= d2^2/d1^2

T2 = T1*( d1^2/d2^2)

T2 = 60 min*(1/7)^2

= 60 min / 49

= 1.22 min --> Answer